1. What is the molarity of a solution that contains 10.0 grams of Silver Nitrate that has been dissolved in 750.0 mL of water? Also, what is the percentage mass, mole fraction of the solute in the solution and mole fraction of the solvent in the solution?
2. You want to create a 0.25 M Potassium Chloride solution. You have mass of 5.00 grams of Potassium Chloride. How much water is needed?
3. What is the molarity of a solution that contains 48 grams of Sodium Chloride and 250 mL of water? What is the molarity ions in the solution?
4. How many mL of hydrogen peroxide are needed to make an 8.50% solution by volume of hydrogen peroxide if you want to make 450.0 mL of solution?
5. What is the molarity of a solution that contains 13.4 grams of calcium chloride dissolved in 655.0 mL of water? What is the molarity ions in the solution?
1
Expert's answer
2020-04-13T02:34:16-0400

1. C = (m / M) / (V / 1000) = (10 / 170) / (750 / 1000) = 0.078 M

m% = 100 × mAgNO3 / msol = 100 ×10 / (750 + 10) = 1.32 %

wAgNO3 = 100 × nAgNO3 / nsol = 100 × (10 / 170) / (10 / 170 + 750 / 18) = 0.15 %

wH2O = 100 - wAgNO3 = 100 - 0.15 = 99.85 %

2. nKCl = 5 / 74.5 = 0.067 mol

VH2O = 1000 × 0.067 / 0.25 = 268 mL

3. CNaCl = (48 / 58.5) / (250 / 1000) = 3.28 M

Every ion molrity is the same, 3.28 M

4. VH2O2 = 450 × (8.5 / 100) = 38.25 mL

5. CCaCl2 = (13.4 / 111) / (655 / 1000) = 0.18 M

CCa2+ = 0.18 M; CCl- = 2 × CCaCl2 = 0.36 M